Which of the following is the equation of the line with slope 7 and y-intercept (0,-6).
a) y=-6x+7
b) -6x+7y=0
c) y=7x-6
d) 7y=-6

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

Which of the following is the equation of the line with slope 7 and y-intercept (0,-6).
a) y=-6x+7
b) -6x+7y=0
c) y=7x-6
d) 7y=-6
---------------
Ans: c
Cheers,
Stan H.

Answer by richwmiller(17219)

y=mx+b
m =slope
b y intercept
so m=7 and y intercept is -6 when x=0 (0,-6)
Which letter is that?
a) slope is -6
b)slope is 6/7
d)slope is 0
that leaves c)slope 7 and y intercept -6